# An Online Learning Approach to Prompt-based Selection of Generative Models and LLMs

### v3

@@ 1 Introduction
Large language models (LLMs) and text-guided generative AI models have found numerous applications in various engineering tasks. A prompt-guided generative AI represents a conditional generative model that produces output samples given an input text prompt. Over the past few years, several frameworks have been developed to train generative models and perform text-guided sample generation in various domains, including text, image, and video (Touvron et al.,, 2023; Gemini-Team et al.,, 2025; OpenAI et al.,, 2024; DeepSeek-AI et al.,, 2025; Rombach et al.,, 2022; Bao et al., 2023b, ; Podell et al.,, 2024; Singer et al.,, 2022; Bai et al.,, 2023; Liu et al.,, 2024). The great number of available LLMs and prompt-guided generative models has led to significant interest in developing evaluation mechanisms to rank the existing models and find the best model from a group of available ones. To address this task, several evaluation metrics have been proposed to quantify the fidelity of samples generated by prompt-based generative models, such as CLIPScore (Hessel et al.,, 2021) and PickScore (Kirstain et al.,, 2023) for prompt-guided image and video generation, and BLEU score (Papineni et al.,, 2002) and BERTscore (Zhang et al.,, 2020) for LLMs.
@@
The existing model selection methodologies commonly aim to identify the generative model with the highest averaged fidelity score over the distribution of text prompts, producing samples that, on average, align the most with input text prompts. A well-known example is the averaged CLIPScore for text-to-image models, measuring the expected alignment between the input text and output image of the model using the CLIP embedding (Radford et al., 2021a, ). While the best-averaged score model selection strategy has been frequently utilized in generative AI applications, this approach does not consider the possibility that the involved models can perform differently across text prompts.
@@
We highlight that the described online learning task can be viewed as a contextual bandit (CB) problem widely studied in the bandit literature (Langford and Zhang,, 2007; Li et al.,, 2010; Chu et al.,, 2011; Agrawal and Goyal,, 2013). In the CB task, the online learner observes a single context variable (the text prompt in our setting) and predicts the best arm for the current input context. Specifically, we focus on the kernel-based method to predict the score of each available model for an incoming prompt. As the text prompt is shared across the models, the CB task may be suboptimally addressed by the common-weight formulation of linear CB (Chu et al.,, 2011) and kernelized CB (Valko et al.,, 2013) algorithms, as these implementations of the algorithms learn a single shared weight vector to predict the reward functions for all the arms (i.e., generative models in our setting). To relax this constraint, we propose Per-Arm Kernelized UCB (PAK-UCB) to address the online prompt-based selection of generative models. According to the PAK-UCB approach, the learner utilizes the computed UCB-scores of arm-specific kernel-based prediction functions to choose the generative model for the incoming text prompt and subsequently update the kernel-based prediction rule based on the generated data for the future rounds. We attempt to theoretically analyze the proposed PAK-UCB and show that its variant could achieve a regret bound of $\widetilde{O}(\sqrt{T})$ over $T$ rounds.
@@
Since the user applying the CB-based model selection approach may have limited compute power and not be able to afford growing computational costs in the online learning process, we propose applying the random Fourier features (RFF) framework (Rahimi and Recht, 2007a, ) to reduce the computational load of PAK-UCB. We discuss that in the proposed PAK-UCB, the computational cost per iteration will grow cubically as $O(t^{3})$ with iteration $t$. We demonstrate that after leveraging the RFF method, our developed proxy RFF-UCB algorithm can effectively approximate the solution to PAK-UCB with the computational costs increasing only linearly $O(t)$.
@@
Finally, we present the results of several numerical experiments to show the efficacy of our proposed PAK-UCB and RFF-UCB in the online selection of prompt-guided generative models and LLMs. In our experiments, we test several pre-trained and simulated text-to-image, image-captioning (image-to-text), and language models, where different models could lead to different rankings across various sample types. Our numerical results suggest a fast convergence of the proposed online learning algorithms to the best prompt-based selection of the available models in response to different prompt types. In our experiments, the proposed PAK-UCB and RFF-UCB could perform better than several baseline methods, including the common-weight implementation of LinUCB (Chu et al.,, 2011) and KernelUCB (Valko et al.,, 2013) methods. The following is a summary of this work’s contributions:
@@ 2 Related Work
(Automatic) Evaluation of conditional generative models. Evaluating the conditional generative models has been studied extensively in the literature. For text-to-image (T2I) generation, earlier methods primarily rely on the Inception score (Salimans et al.,, 2016) and Fréchet inception distance (Heusel et al.,, 2017). More recent works propose reference-free metrics for robust automatic evaluation of T2I and image captioning, with notable examples being CLIPScore (Hessel et al.,, 2021) and PickScore (Kirstain et al.,, 2023). Kim et al., (2022) propose a mutual-information-based metric, which attains consistency across benchmarks, sample parsimony, and robustness. To provide a holistic evaluation of T2I models, several works focus on multi-objective evaluation. Astolfi et al., (2024) propose to evaluate conditional image generation in terms of prompt-sample consistency, sample diversity, and fidelity. Kannen et al., (2024) introduce a framework to evaluate T2I models regarding cultural awareness and cultural diversity. Masrourisaadat et al., (2024) examine the performance of several T2I models in generating images such as human faces and groups and present a social bias analysis. Another line of study explores evaluation approaches using large language models (LLMs). Tan et al., (2024) develop LLM-based evaluation protocols that focus on the faithfulness and text-image alignment. Peng et al., (2024) introduce a GPT-based benchmark for evaluating personalized image generation. In addition, a line of study leverages human feedback for scoring/ranking generated images and improving T2I models (Xu et al.,, 2023). For evaluation of text-to-video (T2V) generation, Huang et al., (2024) introduce VBench as a comprehensive evaluation of T2V models in terms of quality and consistency. Finally, we note the prompt-aware diversity scores, Conditional-Vendi (Jalali et al.,, 2023) and Schur-Complement-Entropy (Ospanov et al., 2024a, ), developed for prompt-guided generative models.
@@
(Kernelized) Contextual bandits.  The contextual bandits (CB) is an efficient framework for online decision-making with side information (Langford and Zhang,, 2007; Foster et al.,, 2018), which is widely adopted in domains such as recommendation system and online advertisement (Li et al.,, 2010) and resource allocations (Lim et al.,, 2024). A key to its formulation is the relationship between the context (vector) and the expected reward. In linear CB, the reward is assumed to be linear to the context vector (Li et al.,, 2010; Chu et al.,, 2011). To incorporate non-linearity, Valko et al., (2013) propose kernelized CB, which assumes the rewards are linear-realizable in a reproducing kernel Hilbert space (RKHS). To address the growing computation, recent work leverages the assumption that the kernel matrix is often approximately low-rank and uses Nyström approximations (Calandriello et al.,, 2019, 2020; Zenati et al.,, 2022).
@@
Prompt-based Selection of LLMs and Generative Models. We note that a line of study proposes offline methods for learning a universal prompt-to-model assignment rule by training a neural network model over a dataset of ranked responses of the models to a large set of prompts (Luo et al.,, 2024; Qin et al.,, 2024; Frick et al.,, 2025). Unlike these methods, our proposed approach follows an online learning strategy which can significantly lower the computational and statistical costs of finding a satisfactory assignment rule. In addition, as we later discuss in the numerical results, the online nature of our proposed algorithm can improve the adaptability of the model to the changes in the user’s prompt distribution and models’ behavior over the iterations of evaluating the models in real time.
@@
Multi-armed bandit selection of generative models. We note that the references (Hu et al., 2025a, ; Rezaei et al.,, 2025) propose multi-armed bandit (MAB) approaches to the selection of unconditional generative models without an input text prompt. Specifically, Hu et al., 2025a propose an online learning framework for selecting generative models and develop the FID-UCB algorithm for the online selection of image-output generative models based on the FID score (Heusel et al.,, 2017). Rezaei et al., (2025) show that a mixture of several unconditional generative models can attain higher diversity scores than each individual model, and propose the Mixture-UCB MAB framework to find the best mixture of the generative models in terms of the kernel distance Bińkowski et al., (2018) and Renyi kernel entropy Jalali et al., (2023); Ospanov et al., 2024b . We note that the mentioned papers study the selection of unconditional generative models, different from our problem setting of selecting the conditional prompt-guided generative models.
@@ 3.1 Kernel Methods and Random Fourier Features
Random Fourier features (RFF). To reduce the computational costs of kernel methods for shift-invariant kernels, Rahimi and Recht, 2007b proposes the random Fourier features framework. Specifically, Bochner’s Theorem (Rudin,, 2017) shows that for every shift-invariant kernel $k(y,y^{\prime})=\kappa(y-y^{\prime})$ satisfying $\kappa(0)=1$, the Fourier transform of $\kappa$ provides a valid probability density function, which we denote by $p\in\Delta({\mathbb{R}}^{d})$, such that $k(y,y^{\prime})=\mathbb{E}_{w\sim p}[e^{iw^{\top}(y-y^{\prime})}],$ with $i$ being the imaginary unit. Following this property, the RFF approach independently samples $w_{1},\cdots,w_{D}\sim p$ and proposes the empirical mean proxy for $k(y,y^{\prime})$ as $\frac{1}{D}\sum_{j=1}^{D}e^{iw_{j}^{\top}(y-y^{\prime})}$. Due to the real Fourier transform of the even function $\kappa$, we can simplify the equation as
@@ 3.2 CLIPScore for Evaluating Text-to-Image Models
CLIPScore (Hessel et al.,, 2021) has been widely-used to evaluate the alignment of samples generated by text-to-image/video (T2I/V) and image captioning models. Let $(y,x)\in{\mathcal{Y}}\times{\mathcal{X}}$ be a text-image pair. We denote by $\bm{c}_{y}\in{\mathbb{S}}^{d-1}:=\{z\in{\mathbb{R}}^{d}:\|z\|_{2}=1\}$ and $\bm{v}_{x}\in{\mathbb{S}}^{d-1}$ the (normalized) embeddings of text $y\in{\mathcal{Y}}$ and image $x\in{\mathcal{X}}$, respectively, both extracted by CLIP (Radford et al., 2021b, ). The CLIPScore (Hessel et al.,, 2021) is given by
@@
4 Prompt-Based Selection as Contextual Bandits
@@ 5 An Optimism-based Approach for Prompt-based Selection
We note that Assumption 1 differs from the standard settings of linear and kernelized contextual bandit (CB) (Chu et al.,, 2011; Valko et al.,, 2013; Zenati et al.,, 2022). Specifically, in the settings of these references, a potentially different context variable is considered for each arm, and the same weight vector is applied across all arms. In the prompt-based model selection setting, the text prompt (i.e., the context variable) is shared among arms (i.e., the generative models) and can result in different performance scores across the models, which cannot be captured by the standard kernelized CB approach. On the other hand, we note that the existing arm-specific bandit algorithms in (Li et al.,, 2010; Xu et al.,, 2020) are designed considering a linear pay-off structure, which is generally different from the kernel-based formulation in our work.
@@ 5.1 The PAK-UCB Algorithm
In this section, we present PAK-UCB in Algorithm 2, an online learning approach to prompt-based model selection. For each incoming prompt, PAK-UCB first estimates the performance scores via kernel ridge regression (KRR) and then picks the model with the highest estimated score. Unlike conventional formulations of LinUCB (Chu et al.,, 2011) and KernelUCB (Valko et al.,, 2013) algorithms, which learn a single reward function shared across all arms, PAK-UCB learns arm-specific functions to predict the score of each model.
@@ 5.2 PAK-UCB with Random Fourier Features
The PAK-UCB solves a KRR for each model at an iteration to estimate the scores, which can be expensive in both computation and memory for a large number of iterations. To address this problem, we leverage the random Fourier features (RFF) sampling (Rahimi and Recht, 2007b, ) for shift-invariant kernel functions, e.g., the RBF (Gaussian) kernel. At a high level, RFF maps the input data, e.g., the prompt (vector) in our setting, to a randomized low-dimensional feature space and then learns a linear model in the resulting random space by solving a standard linear regression problem. Note that following the design of RFF, the inner product of the projected randomized features will be an unbiased estimation of the original kernel inner product.
@@ 6 Numerical Results
Baselines. We compare PAK-UCB-poly3 and RFF-UCB-RBF with three baselines, including 1) KernelUCB-poly3: standard KernelUCB (Valko et al.,, 2013) using the $k_{\textup{poly}}^{1.0}$ kernel, 2) One-arm Oracle: always picking the model with the maximum averaged score, and 3) Random: selecting an model uniformly randomly. In the Appendix, we report results for three additional baselines: 4) PAK-UCB-lin: PAK-UCB with linear kernel, i.e., $k_{\text{lin}}(x_{1},x_{2})=x_{1}^{\top}x_{2}$, which does not incorporate non-linearity in score estimation, 5) LinUCB: standard LinUCB (Chu et al.,, 2011), and 6) Naive-KRR: PAK-UCB-poly3 without exploration, i.e., parameter $\eta=0$, which selects the model with the highest estimated mean conditioned to the prompt.

### v4

@@ 1 Introduction
Large language models (LLMs) and text-guided generative AI models have found numerous applications in various engineering tasks. A prompt-guided generative AI represents a conditional generative model that produces output samples given an input text prompt. Over the past few years, several frameworks have been developed to train generative models and perform text-guided sample generation in various domains, including text, image, and video (Touvron et al., 2023; Gemini-Team et al., 2025; OpenAI et al., 2024; DeepSeek-AI et al., 2025; Rombach et al., 2022; Bao et al., 2023b; Podell et al., 2024; Singer et al., 2022; Bai et al., 2023; Liu et al., 2024). The great number of available LLMs and prompt-guided generative models has led to significant interest in developing evaluation mechanisms to rank the existing models and find the best model from a group of available ones. To address this task, several evaluation metrics have been proposed to quantify the fidelity of samples generated by prompt-based generative models, such as CLIPScore (Hessel et al., 2021) and PickScore (Kirstain et al., 2023) for prompt-guided image and video generation, and BLEU score (Papineni et al., 2002) and BERTscore (Zhang et al., 2020) for LLMs.
@@
The existing model selection methodologies commonly aim to identify the generative model with the highest averaged fidelity score over the distribution of text prompts, producing samples that, on average, align the most with input text prompts. A well-known example is the averaged CLIPScore for text-to-image models, measuring the expected alignment between the input text and output image of the model using the CLIP embedding (Radford et al., 2021b). While the best-averaged score model selection strategy has been frequently utilized in generative AI applications, this approach does not consider the possibility that the involved models can perform differently across text prompts.
@@
We highlight that the described online learning task can be viewed as a contextual bandit (CB) problem widely studied in the bandit literature (Langford & Zhang, 2007; Li et al., 2010; Chu et al., 2011; Agrawal & Goyal, 2013). In the CB task, the online learner observes a single context variable (the text prompt in our setting) and predicts the best arm for the current input context. Specifically, we focus on the kernel-based method to predict the score of each available model for an incoming prompt. As the text prompt is shared across the models, the CB task may be suboptimally addressed by the common-weight formulation of linear CB (Chu et al., 2011) and kernelized CB (Valko et al., 2013) algorithms, as these implementations of the algorithms learn a single shared weight vector to predict the reward functions for all the arms (i.e., generative models in our setting). To relax this constraint, we propose Per-Arm Kernelized UCB (PAK-UCB) to address the online prompt-based selection of generative models. According to the PAK-UCB approach, the learner utilizes the computed UCB-scores of arm-specific kernel-based prediction functions to choose the generative model for the incoming text prompt and subsequently update the kernel-based prediction rule based on the generated data for the future rounds. We attempt to theoretically analyze the proposed PAK-UCB and show that its variant could achieve a regret bound of $\widetilde{O}(\sqrt{T})$ over $T$ rounds.
@@
Since the user applying the CB-based model selection approach may have limited compute power and not be able to afford growing computational costs in the online learning process, we propose applying the random Fourier features (RFF) framework (Rahimi & Recht, 2007b) to reduce the computational load of PAK-UCB. We discuss that in the proposed PAK-UCB, the computational cost per iteration will grow cubically as $O(t^{3})$ with iteration $t$. We demonstrate that after leveraging the RFF method, our developed proxy RFF-UCB algorithm can effectively approximate the solution to PAK-UCB with the computational costs increasing only linearly $O(t)$.
@@
Finally, we present the results of several numerical experiments to show the efficacy of our proposed PAK-UCB and RFF-UCB in the online selection of prompt-guided generative models and LLMs. In our experiments, we test several pre-trained and simulated text-to-image, image-captioning (image-to-text), and language models, where different models could lead to different rankings across various sample types. Our numerical results suggest a fast convergence of the proposed online learning algorithms to the best prompt-based selection of the available models in response to different prompt types. In our experiments, the proposed PAK-UCB and RFF-UCB could perform better than several baseline methods, including the common-weight implementation of LinUCB (Chu et al., 2011) and KernelUCB (Valko et al., 2013) methods. The following is a summary of this work’s contributions:
@@ 2 Related Work
(Automatic) Evaluation of conditional generative models. Evaluating the conditional generative models has been studied extensively in the literature. For text-to-image (T2I) generation, earlier methods primarily rely on the Inception score (Salimans et al., 2016) and Fréchet inception distance (Heusel et al., 2017). More recent works propose reference-free metrics for robust automatic evaluation of T2I and image captioning, with notable examples being CLIPScore (Hessel et al., 2021) and PickScore (Kirstain et al., 2023). Kim et al. (2022) propose a mutual-information-based metric, which attains consistency across benchmarks, sample parsimony, and robustness. To provide a holistic evaluation of T2I models, several works focus on multi-objective evaluation. Astolfi et al. (2024) propose to evaluate conditional image generation in terms of prompt-sample consistency, sample diversity, and fidelity. Kannen et al. (2024) introduce a framework to evaluate T2I models regarding cultural awareness and cultural diversity. Masrourisaadat et al. (2024) examine the performance of several T2I models in generating images such as human faces and groups and present a social bias analysis. Another line of study explores evaluation approaches using large language models (LLMs). Tan et al. (2024) develop LLM-based evaluation protocols that focus on the faithfulness and text-image alignment. Peng et al. (2024) introduce a GPT-based benchmark for evaluating personalized image generation. In addition, a line of study leverages human feedback for scoring/ranking generated images and improving T2I models (Xu et al., 2023). For evaluation of text-to-video (T2V) generation, Huang et al. (2024) introduce VBench as a comprehensive evaluation of T2V models in terms of quality and consistency. Finally, we note the prompt-aware diversity scores, Conditional-Vendi (Jalali et al., 2023) and Schur-Complement-Entropy (Ospanov et al., 2024a), developed for prompt-guided generative models.
@@
(Kernelized) Contextual bandits.  The contextual bandits (CB) is an efficient framework for online decision-making with side information (Langford & Zhang, 2007; Foster et al., 2018), which is widely adopted in domains such as recommendation system and online advertisement (Li et al., 2010) and resource allocations (Lim et al., 2024). A key to its formulation is the relationship between the context (vector) and the expected reward. In linear CB, the reward is assumed to be linear to the context vector (Li et al., 2010; Chu et al., 2011). To incorporate non-linearity, Valko et al. (2013) propose kernelized CB, which assumes the rewards are linear-realizable in a reproducing kernel Hilbert space (RKHS). To address the growing computation, recent work leverages the assumption that the kernel matrix is often approximately low-rank and uses Nyström approximations (Calandriello et al., 2019, 2020; Zenati et al., 2022).
@@
Prompt-based Selection of LLMs and Generative Models. We note that a line of study proposes offline methods for learning a universal prompt-to-model assignment rule by training a neural network model over a dataset of ranked responses of the models to a large set of prompts (Luo et al., 2024; Qin et al., 2024; Frick et al., 2025). Unlike these methods, our proposed approach follows an online learning strategy which can significantly lower the computational and statistical costs of finding a satisfactory assignment rule. In addition, as we later discuss in the numerical results, the online nature of our proposed algorithm can improve the adaptability of the model to the changes in the user’s prompt distribution and models’ behavior over the iterations of evaluating the models in real time.
@@
Multi-armed bandit selection of generative models. We note that the references (Hu et al., 2025a; Rezaei et al., 2025) propose multi-armed bandit (MAB) approaches to the selection of unconditional generative models without an input text prompt. Specifically, Hu et al. (2025a) propose an online learning framework for selecting generative models and develop the FID-UCB algorithm for the online selection of image-output generative models based on the FID score (Heusel et al., 2017). Rezaei et al. (2025) show that a mixture of several unconditional generative models can attain higher diversity scores than each individual model, and propose the Mixture-UCB MAB framework to find the best mixture of the generative models in terms of the kernel distance (Bińkowski et al., 2018) and Renyi kernel entropy (Jalali et al., 2023; Ospanov et al., 2024b). We note that the mentioned papers study the selection of unconditional generative models, different from our problem setting of selecting the conditional prompt-guided generative models.
@@ 3.1 Kernel Methods and Random Fourier Features
Random Fourier features (RFF). To reduce the computational costs of kernel methods for shift-invariant kernels, Rahimi & Recht (2007a) proposes the random Fourier features framework. Specifically, Bochner’s Theorem (Rudin, 2017) shows that for every shift-invariant kernel $k(y,y^{\prime})=\kappa(y-y^{\prime})$ satisfying $\kappa(0)=1$, the Fourier transform of $\kappa$ provides a valid probability density function, which we denote by $p\in\Delta({\mathbb{R}}^{d})$, such that $k(y,y^{\prime})=\mathbb{E}_{w\sim p}[e^{iw^{\top}(y-y^{\prime})}],$ with $i$ being the imaginary unit. Following this property, the RFF approach independently samples $w_{1},\cdots,w_{D}\sim p$ and proposes the empirical mean proxy for $k(y,y^{\prime})$ as $\frac{1}{D}\sum_{j=1}^{D}e^{iw_{j}^{\top}(y-y^{\prime})}$. Due to the real Fourier transform of the even function $\kappa$, we can simplify the equation as
@@ 3.2 CLIPScore for Evaluating Text-to-Image Models
CLIPScore (Hessel et al., 2021) has been widely-used to evaluate the alignment of samples generated by text-to-image/video (T2I/V) and image captioning models. Let $(y,x)\in{\mathcal{Y}}\times{\mathcal{X}}$ be a text-image pair. We denote by $\bm{c}_{y}\in{\mathbb{S}}^{d-1}:=\{z\in{\mathbb{R}}^{d}:\|z\|_{2}=1\}$ and $\bm{v}_{x}\in{\mathbb{S}}^{d-1}$ the (normalized) embeddings of text $y\in{\mathcal{Y}}$ and image $x\in{\mathcal{X}}$, respectively, both extracted by CLIP (Radford et al., 2021a). The CLIPScore (Hessel et al., 2021) is given by
@@
4 Prompt-Based Model Selection as a Contextual Bandit Problem
@@ 5 An Optimism-based Approach for Prompt-based Selection
We note that Assumption 1 differs from the standard settings of linear and kernelized contextual bandit (CB) (Chu et al., 2011; Valko et al., 2013; Zenati et al., 2022). Specifically, in the settings of these references, a potentially different context variable is considered for each arm, and the same weight vector is applied across all arms. In the prompt-based model selection setting, the text prompt (i.e., the context variable) is shared among arms (i.e., the generative models) and can result in different performance scores across the models, which cannot be captured by the standard kernelized CB approach. On the other hand, we note that the existing arm-specific bandit algorithms in (Li et al., 2010; Xu et al., 2020) are designed considering a linear pay-off structure, which is generally different from the kernel-based formulation in our work.
@@ 5.1 The PAK-UCB Algorithm
In this section, we present PAK-UCB in Algorithm 2, an online learning approach to prompt-based model selection. For each incoming prompt, PAK-UCB first estimates the performance scores via kernel ridge regression (KRR) and then picks the model with the highest estimated score. Unlike conventional formulations of LinUCB (Chu et al., 2011) and KernelUCB (Valko et al., 2013) algorithms, which learn a single reward function shared across all arms, PAK-UCB learns arm-specific functions to predict the score of each model.
@@ 5.2 PAK-UCB with Random Fourier Features
The PAK-UCB solves a KRR for each model at an iteration to estimate the scores, which can be expensive in both computation and memory for a large number of iterations. To address this problem, we leverage the random Fourier features (RFF) sampling (Rahimi & Recht, 2007a) for shift-invariant kernel functions, e.g., the RBF (Gaussian) kernel. At a high level, RFF maps the input data, e.g., the prompt (vector) in our setting, to a randomized low-dimensional feature space and then learns a linear model in the resulting random space by solving a standard linear regression problem. Note that following the design of RFF, the inner product of the projected randomized features will be an unbiased estimation of the original kernel inner product.
@@ 6 Numerical Results
Baselines. We compare PAK-UCB-poly3 and RFF-UCB-RBF with three baselines, including 1) KernelUCB-poly3: standard KernelUCB (Valko et al., 2013) using the $k_{\textup{poly}}^{1.0}$ kernel, 2) One-arm Oracle: always picking the model with the maximum averaged score, and 3) Random: selecting an model uniformly randomly. In the Appendix, we report results for three additional baselines: 4) PAK-UCB-lin: PAK-UCB with linear kernel, i.e., $k_{\text{lin}}(x_{1},x_{2})=x_{1}^{\top}x_{2}$, which does not incorporate non-linearity in score estimation, 5) LinUCB: standard LinUCB (Chu et al., 2011), and 6) Naive-KRR: PAK-UCB-poly3 without exploration, i.e., parameter $\eta=0$, which selects the model with the highest estimated mean conditioned to the prompt.